The measure of one angle is 15 degrees more than half the measure of its complement. Find the measure of the two angles.
Answer #1
A = Compliment = X Deg.
B = (X/2 + 15) Deg.

The sum of complimentary angles = 90 Deg.
X + (X/2+15) = 90 Deg.
Multiply both sides by 2:
2x + x + 30 = 180.
3x = 180 - 30 = 150.

X = 150 / 3 = 50 Deg.
x/2 + 15 = 50/2 + 15 = 40 Deg.
answered by: Sage